An electronic company in Laguna manufactures resistors that have a mean resistance of 120 ohms and a standard deviation of 13 ohms. Find the probability that a random sample of 40 resistors will have an average resistance greater than 114 ohms.


1
Expert's answer
2022-03-03T09:49:23-0500

Let "X=" an average resistance: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=120\Omega, \sigma=13\Omega , n=40."


"P(X>114)=P(Z>\dfrac{114-120}{13/\sqrt{40}})"

"\approx P(Z>-2.9190)\approx0.9982"